OpenStudy (anonymous):

If g(x) = 3x + 5, what is g–1(x)?

OpenStudy (saifoo.khan):

g inverse right?

OpenStudy (saifoo.khan):

g(x) = 3x + 5 y = 3x + 5 Make x the subject: (y -5)/3 = x

OpenStudy (saifoo.khan):

Final step: replace y with x. g-1(x) = (x-5)/3

OpenStudy (anonymous):

So the final answeris?

OpenStudy (saifoo.khan):

What do you think? :D

OpenStudy (anonymous):

x-3/5?